# YourMT3+: Multi-instrument Music Transcription with Enhanced Transformer Architectures and Cross-dataset Stem Augmentation

### v1

@@ Abstract
Multi-instrument music transcription aims to convert polyphonic music recordings into musical scores assigned to each instrument. This task is challenging for modeling as it requires simultaneously identifying multiple instruments and transcribing their pitch and precise timing, and the lack of fully annotated data adds to the training difficulties. This paper introduces YourMT3+, a suite of models for enhanced multi-instrument music transcription based on the recent language token decoding approach of MT3. We strengthen its encoder by adopting a hierarchical attention transformer in the time-frequency domain and integrating a mixture of experts (MoE). To address data limitations, we introduce a new multi-channel decoding method for training with incomplete annotations and propose intra- and cross-stem augmentation for dataset mixing. Our experiments demonstrate direct vocal transcription capabilities, eliminating the need for voice separation pre-processors. Benchmarks across ten public datasets show our models’ competitiveness with, or superiority to, existing transcription models. Further testing on pop music recordings highlights the limitations of current models. Fully reproducible code and datasets are available at https://github.com/mimbres/YourMT3.
@@ 1 Introduction
The key challenge of this research is multi-instrument AMT: identification and transcription of various instruments with vocals from music recordings. Recently, there has been notable progress in this field: MT3 [4] utilized a MIDI-like decoding transformer, while PerceiverTF [5] employed a spectral attention transformer that generates conventional piano-roll. Unfortunately, the absence of fully reproducible code for these models has been a significant limitation for replication and further research. Our replication of MT3, trainable from scratch, is dubbed as YourMT3 [6]. Based on this, we propose YourMT3+, a hybrid architecture that incorporates advanced architectures and training methods for further enhancements. YourMT3+ and its variants differ from prior works [4, 5] in the following key aspects:
  • Encoder: PerceiverTF [5], which generated piano-rolls, is now trained with the MT3 framework to generate note event tokens. We replaced MT3’s encoder with PerceiverTF featuring spectral cross attention (SCA). Additionally, replacing its feedforward network (FFN) with a mixture of experts (MoE) [7], denoted as YPTF.MoE, demonstrates promising results.
  • Decoder: Tokens for singing transcription have been further defined. We introduce a multi-channel decoder that replaces MT3’s single-channel decoder [4]. This enables task-query based training and the use of partially annotated data, improving performance.
@@ 3 Model
In the YourMT3+ model taxonomy, we designate models with the same architecture and training method as MT3 as YMT3. Models that replace the encoder with PerceiverTF, trained using MT3’s single-channel decoder and our stem augmentation, are termed YPTF+Single. The variant replacing the encoder’s FFN with a mixture of experts (MOE) is called YPTF.MoE. These models have demonstrated that the PerceiverTF encoder can efficiently process language-style tokens, such as MIDI, instead of using piano-roll. Additionally, we propose a multi-channel decoder that assigns instrument groups per channel and masks loss for unannotated instruments, allowing training with incomplete labels. The final model, which integrates all these features, is termed YPTF.MoE+Multi.
@@ 3.2 Encoder
MoE: Our models, designated as YPTF.MoE, replace the FFNs in latent and temporal transformer blocks with mixture of experts (MoE) layers from [7], routing attention output to two out of eight experts (see Supplemental B.5). In our experiments, MoE increased the model complexity by about 5% while improving performance across various datasets. Unlike PerceiverTF, we use RoPE [17] in every sub-block of the encoder to integrate positional information through rotation matrices, replacing trainable position embedding (PE), and pre-LayerNorm with pre-RMSNorm. However, these modifications only offered minor benefits in memory and computation without significantly impacting performance.
@@ 4.1 Intra-stem Augmentation
with $i\in\{1,2,...,N\}$ where $N>1$. Here, $p$ represents the probability of each stem being selected, set to 0.7 as default. Each $x_{i}$ may or may not be chosen with a probability $p$, resulting in $\hat{X}_{\text{in}}$ containing various combinations of stems from the original segment $X$.
@@ 4.3 Pitch-shifting
During training, pitch-shifting is applied to a batch, after applying cross-dataset stem augmentation. The phase vocoder-based pitch-shifting on the GPU was adapted from the TorchAudio [^1] https://pytorch.org/audio library. We used the default settings with the exception of setting $\textit{nFFT}=512$ for time-stretching. All the elements within the batch are randomly assigned to one of five groups, each with an equal chance of selection. Each group is then subjected to a pitch shift by an interval of -2, -1, 0, +1, or +2 semitones.
@@ 5.1 Experimental Setup
Data Preparation: Table 1 lists the datasets used for training and evaluating our model. We offer a software package, [an anonymised link], for dataset setup and split information to ensure reproducibility of our results. Audio data was converted into 16 kHz mono WAV format. Stems were stored as arrays, and mix-tracks as WAV files, also treating stemless tracks as mix-tracks. For training our Single decoder models on MIR-ST500 [19] and CMedia [20] , we produced singing and accompaniment stems using a pre-trained separation model [13]. With the Multi decoder, we also incorporated the original mix tracks from these datasets.
@@ 5.2 Results and Discussion
Performance on Pop Music: As seen at the bottom of Table 3, our model demonstrated competitive performance on the synthetic dataset [29] compared to other multi-AMT models. In Figure 3, our final model achieved 50 to 90% performance for most instruments, except for a few non-mainstream ones like chromatic percussion (c. perc) and synth pad (s.pad) in the synthetic dataset. However, it significantly underperformed on commercial pop music recordings as shown in Figure 3. Particularly for non-main instruments (excluding piano, bass, vocals, and drums), our models performed below 10%. This indicates that relying solely on synthetic datasets is insufficient for modeling the diverse timbres of pop music. Furthermore, except for the piano, all the pitched instruments showed a significant gap in the chroma-level metric, suggesting substantial octave errors and hinting that more varied pitch-shifting could be beneficial.

### v2

@@ Abstract
Multi-instrument music transcription aims to convert polyphonic music recordings into musical scores assigned to each instrument. This task is challenging for modeling as it requires simultaneously identifying multiple instruments and transcribing their pitch and precise timing, and the lack of fully annotated data adds to the training difficulties. This paper introduces YourMT3+, a suite of models for enhanced multi-instrument music transcription based on the recent language token decoding approach of MT3. We enhance its encoder by adopting a hierarchical attention transformer in the time-frequency domain and integrating a mixture of experts. To address data limitations, we introduce a new multi-channel decoding method for training with incomplete annotations and propose intra- and cross-stem augmentation for dataset mixing. Our experiments demonstrate direct vocal transcription capabilities, eliminating the need for voice separation pre-processors. Benchmarks across ten public datasets show our models’ competitiveness with, or superiority to, existing transcription models. Further testing on pop music recordings highlights the limitations of current models. Fully reproducible code and datasets are available with demos at https://github.com/mimbres/YourMT3.
@@ 1 Introduction
The key challenge of this research is multi-instrument AMT: identification and transcription of various instruments with vocals from music recordings. Recently, there has been notable progress in this field: MT3 [4] utilized a MIDI-like decoding transformer, while PerceiverTF [5] employed a spectral attention transformer that generates conventional piano-roll. Unfortunately, the absence of fully reproducible code for these models has been a significant limitation for replication and further research. Our replication of MT3, trainable from scratch, is dubbed as YourMT3 [6]. Based on this, we propose YourMT3+, a hybrid architecture that incorporates advanced architectures and training methods for further enhancements. YourMT3+ and its variants differ from prior work [4, 5] in the following key aspects:
  • Enhanced Encoder: PerceiverTF [5], which generated piano-rolls, is now trained with the MT3 framework to generate note event tokens. We replaced MT3’s encoder with PerceiverTF featuring spectral cross attention (SCA). Additionally, replacing its feedforward network (FFN) with a mixture of experts (MoE) [7], denoted as YPTF.MoE, demonstrates promising results.
  • Multi-channel Decoder: In addition to General MIDI tokens, singing transcription tokens have been further defined. We introduce a multi-channel decoder that replaces MT3’s single-channel decoder [4]. This enables task-query based training and the use of partially annotated data, improving performance.
@@ 3 Model
In the YourMT3+ taxonomy, YMT3 models match MT3’s [4] architecture and training. YPTF+Single models use PerceiverTF (PTF) encoder with MT3’s single-channel decoder and stem augmentation. Our empirical finding demonstrates that PTF’s hierarchical attention with instrument-group sub-task queries enhances multi-instrument AMT in complex mixtures. YPTF.MoE replaces the encoder’s FFN with mixture of experts (MoE), enabling task-specific encodings in multi-dataset training. These models efficiently process MIDI tokens instead of piano-roll. Our multi-channel decoder assigns instrument groups per channel and masks loss for unannotated instruments, allowing training with incomplete labels. The final YPTF.MoE+Multi model integrates all these features.
@@ 3.2 Encoder
MoE: YPTF.MoE models replace FFNs in latent and temporal transformer blocks with MoE layers [7], routing attention to two of eight experts. Using two experts gave better results than one or four; see Supplemental B.5. In our experiments, MoE increased the model complexity by about 5% while improving performance across various datasets. Unlike PerceiverTF, we use RoPE [17] in every sub-block of the encoder to integrate positional information through rotation matrices, replacing trainable position embedding (PE), and pre-LayerNorm with pre-RMSNorm. However, these modifications only offered minor benefits in memory and computation without significantly impacting performance.
@@ 4.1 Intra-stem Augmentation
with $i\in\{1,2,...,N\}$ where $N>1$. Here, $p$=$0.7$ by default, is the probability of each stem being selected. Each $x_{i}$ is chosen with $p$, creating $\hat{X}_{\text{in}}$ with various combinations of stems from $X$. A larger $p$ increases active stems and task difficulty. The sweet spot was between 0.6 and 0.8, increasing with model size and training time.
@@ 4.3 Pitch-shifting
We apply GPU-based phase vocoder pitch-shifting adapted from TorchAudio [^1] https://pytorch.org/audio after cross-dataset stem augmentation, using default settings except nFFT=512 for time-stretching. Batch elements are randomly assigned to five groups, each shifted by -2, -1, 0, +1, or +2 semitones. Notably, as will be discussed in Section 5.2, pitch shifting’s inconsistent benefits across datasets were resolved by MoE models’ increased capacity.
@@ 5.1 Experimental Setup
Data Preparation: Table 1 lists the datasets used for training and evaluating our model. We offer a software package for dataset setup and split information to ensure reproducibility of our results. Audio data was converted into 16 kHz mono WAV format. Stems were stored as arrays, and mix-tracks as WAV files, also treating stemless tracks as mix-tracks. For training our Single decoder models on MIR-ST500 [19] and CMedia [20] , we produced singing and accompaniment stems using a pre-trained separation model [13]. With the Multi decoder, we also incorporated the original mix tracks from these datasets.
@@ 5.2 Results and Discussion
Performance on Pop Music: As seen at the bottom of Table 3, our model demonstrated competitive performance on the synthetic dataset [29] compared to other multi-AMT models. In Figure 3, our final model achieved 50 to 90% performance for most instruments, except for a few non-mainstream ones like chromatic percussion (c. perc) and synth pad (s.pad) in the synthetic dataset. However, a significant limitation emerged in its performance on commercial pop music recordings, as shown in Figure 3. Particularly for non-main instruments (excluding piano, bass, vocals, and drums), our models performed below 10%. This suggests potential biases introduced by training primarily on synthetic datasets, which may not fully cover the diverse timbres of pop music. Furthermore, except for the piano, all the pitched instruments showed a significant gap in the chroma-level metric, suggesting substantial octave errors and hinting that more varied pitch-shifting could be beneficial.